The position of a particle as a function of time i
Answer #1

Assume that the particle is travelling along a straight line.

Q3.

At time t_1=0.6\,s position of particle is x_1=1.2\,m. At time t_2=7.2\,s  position of particle is x_2=2.8\,m

Average velocity of particle between t_1=0.6\,s and t_2=7.2\,s is V_{avg}=\frac{\text{displacment}}{\text{time taken}}=\frac{x_2-x_1}{t_2-t_1}=\frac{2.8-1.2}{7.2-0.6}=0.24\,m/s

Q4.

Average speed of particle between t_1=0.6\,s and t_2=7.2\,s is |\vec{V}|_{avg}=\frac{\text{distance travelled}}{\text{time taken}}=\frac{17.6}{6.6}=2.7\,m/s

Q5.

At time t_1=0.6\,s instantaneous velocity of particle is v_1=2\,m/s. At time t_2=7.2\,s velocity of particle is v_2=4\,m/s

Average acceleration of particle between t_1=0.6\,s and t_2=7.2\,s is a_{avg}=\frac{v_2-v_1}{t_2-t_1}=\frac{4-2}{7.2-0.6}=0.30\,m/s^2